If f(x)= 4x-2, then f(x+h)-f(x)=
Found 2 solutions by nerdybill, Fombitz:
Answer by nerdybill(7384)

If f(x)= 4x-2, then f(x+h)-f(x)=
.
Plug in the given values:
f(x+h) = 4(x+h) - 2
and, of course:
f(x) = 4x -2
.
So,
f(x+h)-f(x)= 4(x+h) - 2 - (4x -2)
f(x+h)-f(x)= 4x+4h-2 - (4x -2)
f(x+h)-f(x)= 4x+4h-2 -4x+2
f(x+h)-f(x)= 4x+4h-4x
f(x+h)-f(x)= 4h

Answer by Fombitz(32388)